# Robust Max Statistics for High-Dimensional Inference

### v1

@@ Abstract
Although much progress has been made in the theory and application of bootstrap approximations for max statistics in high dimensions, the literature has largely been restricted to cases involving light-tailed data. To address this issue, we propose an approach to inference based on robust max statistics, and we show that their distributions can be accurately approximated via bootstrapping when the data are both high-dimensional and heavy-tailed. In particular, the data are assumed to satisfy an extended version of the well-established $L^{4}$-$L^{2}$ moment equivalence condition, as well as a weak variance decay condition. In this setting, we show that near-parametric rates of bootstrap approximation can be achieved in the Kolmogorov metric, independently of the data dimension. Moreover, this theoretical result is complemented by favorable empirical results involving both synthetic data and an application to financial data.
@@ 1 Introduction
Over the past decade, distributional approximation results for max statistics have become a prominent topic in high-dimensional inference. A prototypical example of such a statistic has the form $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$, where $\bar{X}\in\mathbb{R}^{p}$ is the sample mean vector of $n$ observations and $\mu=\mathbf{E}(\bar{X})$, but numerous variants arise in diverse contexts. Indeed, one of the main drivers of research on this topic is that many high-dimensional inference tasks can be unified within the problem of approximating the distribution of $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$, or some adaptation of it. For instance, such approximations can be directly applied to construct simultaneous tests and confidence intervals for coordinate-wise means $\mu_{1},\dots,\mu_{p}$. More broadly, other applications include detection of treatment effects (Sun:2022, ), error estimation for sample covariance matrices (Lopes:Bernoulli, ), post-selection inference (Kuchibhotla:2020:PSI, ), change-point detection (Yu:2021, ), confidence bands in non-parametric regression (Singh:2023, ), tests for shape restrictions (Chetverikov:shape, ), and more. Meanwhile, another major reason why max statistics have attracted growing interest is that bootstrap methods can accurately approximate their distributions when $p$ is much larger than the sample size $n$, which has been demonstrated by a cascade of theoretical advances CCK:2013 ; CCK:2017 ; Deng:2020 ; Kuchibhotla2020high ; lopes2020bootstrapping ; Kuchibhota_2021 ; Lopes:2022:AOS ; CCK:2023 ; Fang:2023 ; Koike:2024 .
@@
Despite the substantial innovations that have been made in bootstrap approximations for max statistics, there is an Achilles heel that continues to hinder much of the research in this area. Namely, there is a widespread reliance on the assumption that the covariates have light tails, e.g. sub-Gaussian or sub-exponential. Moreover, there are empirical and theoretical results suggesting that light tails are necessary for bootstrap methods to successfully approximate the distributions of conventional max statistics in high dimensions (zhang2017gaussian, ; Giessing:2020, ; Preinerstorfer:2024, ). For instance, the simulations in (Giessing:2020, ) show that the Gaussian multiplier bootstrap performs poorly for $\max_{1\leq j\leq p}\sqrt{n}|\bar{X}_{j}-\mu_{j}|$ when the covariates have heavy tails and $p\gg n$. From a theoretical standpoint, it has also been proven that there is a moment-dependent phase transition governing the success of Gaussian approximations for $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$ (zhang2017gaussian, ; Preinerstorfer:2024, ). That is, if $W\in\mathbb{R}^{p}$ is a centered Gaussian vector having the same covariance matrix as $\sqrt{n}(\bar{X}-\mu)$, then the Kolmogorov distance between $\max_{1\leq j\leq p}W_{j}$ and $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$ will approach 0 or 1 in the limit that $n$ and $p$ jointly diverge, depending on whether the covariates have enough moments. This breakdown of Gaussian approximations suggests that similar behavior should occur for bootstrap approximations—especially in the case of the Gaussian multiplier bootstrap, which seeks to mimic the behavior of $\sqrt{n}(\bar{X}-\mu)$ by generating random vectors from a centered Gaussian distribution whose covariance matrix is an estimate for that of $\sqrt{n}(\bar{X}-\mu)$.
@@
Due to the issues just mentioned, there are strong motivations to extend bootstrap methods involving max statistics so that they can be applied reliably to high-dimensional data with heavy tails. However, the research in this direction is still at a very early stage, and there are just a couple of previous works that have given it attention. The first of these works briefly outlined an approach that combines truncation with permutation-based sampling (Lou:2017, ), but it was ultimately not pursued as a practical method for inference. More recently, the state-of-the-art paper (Fan:2023, ) proposed a weighted bootstrap for a max statistic of the form $\max_{1\leq j\leq p}\sqrt{n}|\hat{\theta}_{j}-\theta_{j}|$, where $\theta_{j}$ denotes the so-called “pseudomedian” of the $j$th covariate, and $\hat{\theta}_{j}$ is the classical Hodges-Lehmann estimator for $\theta_{j}$ (Hodges:1963, ).
@@
While the approach in (Fan:2023, ) achieved major progress by delivering robust simultaneous inference for $\theta_{1},\dots,\theta_{p}$, it still has some essential limitations. One is that the pseudomedians can be unsatisfactory substitutes for the means $\mu_{1},\dots,\mu_{p}$, particularly in cases of asymmetric distributions, for which $\theta_{j}$ and $\mu_{j}$ may be quite different. A related issue is that an approach based on pseudomedians does not extend naturally to suprema of zero-mean empirical processes, which appear frequently in applications of bootstrap approximations for max statistics (CCK_bands, ; Chen:ridges, ; Han:Bernoulli, ; Chen:Jackknife, ; Dette:Banach, ; Lopes:Bernoulli, ; Giessing:processes, ). Another issue is that the method in (Fan:2023, ) produces simultaneous confidence intervals for $\theta_{1},\dots,\theta_{p}$ that are only theoretically justified when they all have the same width, which is impractical if the covariates fluctuate over different scales. Lastly, the available theoretical analysis for $\max_{1\leq j\leq p}\sqrt{n}|\hat{\theta}_{j}-\theta_{j}|$ establishes a near $n^{-1/4}$ rate of bootstrap approximation in the Kolmogorov metric, which does not align with other recent results for max statistics that establish near $n^{-1/2}$ rates in the setting of light-tailed data lopes2020bootstrapping ; Lopes:2022:AOS ; CCK:2023 ; Fang:2023 ; Koike:2024 .
@@
In the current paper, we propose to bootstrap a robust max statistic that enables simultaneous inference on the means $\mu_{1},\dots,\mu_{p}$ and overcomes the difficulties described above. Our approach is designed in terms of three ingredients: truncation, partial standardization, and the median-of-means (mom) technique  (Nemirovsky1983, ; lugosi2019mean, ). To briefly lay out the main ideas, let $X_{1},\dots,X_{p}\in\mathbb{R}^{p}$ be i.i.d. observations with $\textup{var}(X_{1j})=\sigma_{j}^{2}$, and $\mu_{j}=\mathbf{E}(X_{1j})$ as before. Also, let $\hat{\sigma}_{1}^{2},\dots,\hat{\sigma}_{p}^{2}$ denote variance estimates that will be constructed from a small hold-out set via mom, and define the truncation function $\varphi_{t}(x)=\textup{sgn}(x)(|x|\wedge t)$ for any $x\in\mathbb{R}$ and $t\geq 0$, where $a\wedge b=\min\{a,b\}$. In this notation, the proposed robust max statistic is defined by
@@
With regard to theoretical analysis, we focus on a setting where the tails of the data are quantified by a variant of the $L^{4}$-$L^{2}$ moment equivalence condition, which has gained increasing currency in the high-dimensional robustness literature lugosi2019mean ; ke2019user ; mendelson2020robust ; roy2021empirical ; abdalla2022covariance . Specifically, we assume there is some $\delta>0$ such that the bound $\|\langle v,X_{1}-\mu\rangle\|_{L^{4+\delta}}\lesssim\|\langle v,X_{1}-\mu\rangle\|_{L^{2}}$ holds for all $v\in\mathbb{R}^{p}$, and in Proposition 1, we show that this condition is satisfied by heavy-tailed instances of well-known models. The other primary structural assumption in our analysis is that the covariates have a weak variance decay property of the form $\sigma_{(j)}^{2}\asymp j^{-2\beta}$ for some fixed $\beta>0$, where $\sigma_{(1)}^{2}\geq\cdots\geq\sigma_{(p)}^{2}$ are the sorted coordinate-wise variances. Notably, the decay is referred to as weak because the parameter $\beta$ is allowed to be arbitrarily small. Furthermore, it is known that this type of structure arises naturally in a variety of high-dimensional contexts that are related to principal components analysis and functional data analysis, among others (lopes2020bootstrapping, ). Under the complete set of conditions given in Assumption 1, our main result shows that with high probability, the Kolmogorov distance $\sup_{s\in\mathbb{R}}|\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)|$ is nearly of order $n^{-1/2}$, where $\mathbf{P}(\cdot|X)$ denotes probability that is conditional on all of the observations.
From a practical standpoint, the proposed method has several strengths. First, the method does not require fine tuning, which is demonstrated by the fact that we use the simple choices of $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ and $\tau=0.9$ throughout all of the experiments presented in Section 4. Second, we show that the method reliably produces well-calibrated tests and confidence intervals across many conditions—including heavy-tailed data generated from separable and elliptical distributions, as well heavy-tailed functional data with rough sample paths. Third, the simulation results reveal that the proposed method performs favorably in comparison to the pseudomedian approach in (Fan:2023, ). Lastly, we present a financial application involving the detection of stocks that are suitable for “pairs trading”, which shows that the proposed method typically yields smaller $p$-values than those obtained from the pseudomedian approach.
@@ 2 Method
Here, we provide the details for constructing the bootstrap sample $\mathcal{M}_{n}^{*}$, as well as simultaneous confidence intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for the coordinate-wise means $\mu_{1},\dots,\mu_{p}$. Further applications of these intervals to various testing problems will be covered later in Section 4.
@@
Next, we turn to the construction of simultaneous confidence intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\mu_{1},\dots,\mu_{p}$. Let $1-\alpha$ denote the nominal simultaneous coverage probability, and let $\hat{q}_{+}(1-\alpha/2)$ denote the empirical $(1-\alpha/2)$-quantile of a collection of bootstrap samples generated in the manner of $\mathcal{M}_{n}^{*}$. Also, let ¯$\underline{\mathcal{M}}_{n}^{*}$ denote the counterpart of $\mathcal{M}_{n}^{*}$ that is obtained by replacing $\max_{1\leq j\leq p}$ with $\min_{1\leq j\leq p}$ in equation (2), and let $\hat{q}_{-}(\alpha/2)$ denote the empirical $(\alpha/2)$-quantile of a collection of bootstrap samples generated in the manner of ¯$\underline{\mathcal{M}}_{n}^{*}$. In this notation, the confidence interval ^$\hat{\mathcal{I}}_{j}$ is defined by
^$\hat{\mathcal{I}}_{j}=\bigg{\{}x\in\mathbb{R}\,:\,\hat{q}_{-}(\alpha/2)\leq\frac{1}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}\sum_{i=1}^{n}\varphi_{\hat{t}_{j}}(X_{ij}-x)\leq\hat{q}_{+}(1-\alpha/2)\bigg{\}}.$ (7)
Due to the fact that the functions $\varphi_{\hat{t}_{1}}(\cdot),\dots,\varphi_{\hat{t}_{p}}(\cdot)$ are monotone, it is straightforward to compute all the endpoints of ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$.
To comment on the role of the partial standardization parameter $\tau\in[0,1]$, it provides a way to balance two opposing effects that occur in the extreme cases when $\tau$ is equal to 0 or 1. When $\tau=0$, all of the intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ have the same width, which is clearly undesirable when the covariates fluctuate over different scales. Alternatively, when $\tau=1$, all of the covariates will be on approximately “equal footing”, which will tend to make the max statistic $\mathcal{M}_{n}$ sensitive to all $p$ dimensions. This is undesirable in high-dimensional situations where the covariates fluctuate over different scales, because it eliminates a form of low-dimensional structure that can simplify the behavior of $\mathcal{M}_{n}$ when $\tau<1$. To see this, consider a case where $\tau=0$ and $\sigma_{1},\dots,\sigma_{d}$ are much larger than $\sigma_{d+1},\dots,\sigma_{p}$ for some $d\ll p$. In this case, the maximizing index for $\mathcal{M}_{n}$ is likely to reside in the small subset $\{1,\dots,d\}\subset\{1,\dots,p\}$. Thus, the behavior of $\mathcal{M}_{n}$ will be mainly governed by the first $d$ covariates, which intuitively reduces the effective dimension of the problem of approximating the distribution of $\mathcal{M}_{n}$. Accordingly, as was originally proposed in (lopes2020bootstrapping, ), it is natural to select an intermediate value of $\tau$ between 0 and 1 that can mitigate the unwanted effects that occur at $\tau\in\{0,1\}$.

### v2

@@ Abstract
Although much progress has been made in the theory and application of bootstrap approximations for max statistics in high dimensions, the literature has largely been restricted to cases involving light-tailed data. To address this issue, we propose an approach to inference based on robust max statistics, and we show that their distributions can be accurately approximated via bootstrapping when the data are both high-dimensional and heavy-tailed. In particular, the data are assumed to satisfy an extended version of the well-established $L^{4}$-$L^{2}$ moment equivalence condition, as well as a weak variance decay condition. In this setting, we show that near-parametric rates of bootstrap approximation can be achieved in the Kolmogorov metric, independently of the data dimension. Moreover, this theoretical result is complemented by encouraging empirical results involving both Euclidean and functional data.
@@ 1 Introduction
Over the past decade, distributional approximation results for max statistics have become a prominent topic in high-dimensional inference. A prototypical example of such a statistic has the form $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$, where $\bar{X}\in\mathbb{R}^{p}$ is the sample mean vector of $n$ observations and $\mu=\mathbf{E}(\bar{X})$, but numerous variants arise in diverse contexts. Indeed, one of the main drivers of research on this topic is that many high-dimensional inference tasks can be unified within the problem of approximating the distribution of $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$, or some adaptation of it. For instance, such approximations can be directly applied to construct simultaneous tests and confidence intervals for coordinate-wise means $\mu_{1},\dots,\mu_{p}$. More broadly, other applications include detection of treatment effects (Sun:2022), error estimation for sample covariance matrices (Lopes:Bernoulli), post-selection inference (Kuchibhotla:2020:PSI), change-point detection (Yu:2021), confidence bands in non-parametric regression (Singh:2023), tests for shape restrictions (Chetverikov:shape), and more. Meanwhile, another major reason why max statistics have attracted growing interest is that bootstrap methods can accurately approximate their distributions when $p$ is much larger than the sample size $n$, which has been demonstrated by a cascade of theoretical advances CCK:2013; CCK:2017; Deng:2020; Kuchibhotla2020high; lopes2020bootstrapping; Kuchibhota_2021; Lopes:2022:AOS; CCK:2023; Fang:2023; Koike:2024.
@@
Despite the substantial innovations that have been made in bootstrap approximations for max statistics, there is an Achilles heel that continues to hinder much of the research in this area. Namely, there is a widespread reliance on the assumption that the covariates have light tails, e.g., sub-Gaussian or sub-exponential. Moreover, there are empirical and theoretical results suggesting that light tails are necessary for bootstrap methods to successfully approximate the distributions of conventional max statistics in high dimensions (zhang2017gaussian; Giessing:2020; Preinerstorfer:2024). For instance, the simulations in (Giessing:2020) show that the Gaussian multiplier bootstrap performs poorly for $\max_{1\leq j\leq p}\sqrt{n}|\bar{X}_{j}-\mu_{j}|$ when the covariates have heavy tails and $p\gg n$. From a theoretical standpoint, it has also been proven that there is a moment-dependent phase transition governing the success of Gaussian approximations for $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$ (zhang2017gaussian; Preinerstorfer:2024). That is, if $W\in\mathbb{R}^{p}$ is a centered Gaussian vector having the same covariance matrix as $\sqrt{n}(\bar{X}-\mu)$, then the Kolmogorov distance between $\max_{1\leq j\leq p}W_{j}$ and $\max_{1\leq j\leq p}\sqrt{n}(\bar{X}_{j}-\mu_{j})$ may or may not vanish in the limit that $n$ and $p$ jointly diverge, depending on whether the covariates have enough moments. This breakdown of Gaussian approximations suggests that similar behavior should occur for bootstrap approximations—especially in the case of the Gaussian multiplier bootstrap, which seeks to mimic the behavior of $\sqrt{n}(\bar{X}-\mu)$ by generating random vectors from a centered Gaussian distribution whose covariance matrix is an estimate for that of $\sqrt{n}(\bar{X}-\mu)$.
@@
Due to the issues just mentioned, there are strong motivations to extend bootstrap methods involving max statistics so that they can be applied reliably to high-dimensional data with heavy tails. However, the research in this direction is still at a very early stage, and there are just a couple of previous works that have given it attention. The first of these works briefly outlined an approach that combines truncation with permutation-based sampling (Lou:2017), but it was ultimately not pursued as a practical method for inference. More recently, the state-of-the-art paper (Fan:2023) proposed a weighted bootstrap for a max statistic of the form $\max_{1\leq j\leq p}\sqrt{n}|\hat{\theta}_{j}-\theta_{j}|$, where $\theta_{j}$ denotes the so-called “pseudomedian” of the $j$th covariate, and $\hat{\theta}_{j}$ is the classical Hodges-Lehmann estimator for $\theta_{j}$ (Hodges:1963).
@@
While the approach in (Fan:2023) achieved major progress by delivering robust simultaneous inference for $\theta_{1},\dots,\theta_{p}$, it still has some essential limitations. One is that the pseudomedians can be unsatisfactory substitutes for the means $\mu_{1},\dots,\mu_{p}$, particularly in cases of asymmetric distributions, for which $\theta_{j}$ and $\mu_{j}$ may be quite different. A related issue is that an approach based on pseudomedians does not extend naturally to suprema of zero-mean empirical processes, which appear frequently in applications of bootstrap approximations for max statistics (CCK_bands; Chen:ridges; Han:Bernoulli; Chen:Jackknife; Dette:Banach; Lopes:Bernoulli; Giessing:processes). Another issue is that the method in (Fan:2023) produces simultaneous confidence intervals for $\theta_{1},\dots,\theta_{p}$ that are only theoretically justified when they all have the same width, which is impractical if the covariates fluctuate over different scales. Lastly, the available theoretical analysis for $\max_{1\leq j\leq p}\sqrt{n}|\hat{\theta}_{j}-\theta_{j}|$ establishes a near $n^{-1/4}$ rate of bootstrap approximation in the Kolmogorov metric, which does not align with other recent results for max statistics that establish near $n^{-1/2}$ rates in the setting of light-tailed data lopes2020bootstrapping; Lopes:2022:AOS; CCK:2023; Fang:2023; Koike:2024.
@@
In the current paper, we propose to bootstrap a robust max statistic that enables simultaneous inference on the means $\mu_{1},\dots,\mu_{p}$ and overcomes the difficulties described above. Our approach is designed in terms of three ingredients: truncation, partial standardization, and the median-of-means (mom) technique  (Nemirovsky1983; lugosi2019mean). To briefly lay out the main ideas, let $X_{1},\dots,X_{n}\in\mathbb{R}^{p}$ be i.i.d. observations with $\textup{var}(X_{1j})=\sigma_{j}^{2}$, and $\mu_{j}=\mathbf{E}(X_{1j})$ as before. Also, let $\hat{\sigma}_{1}^{2},\dots,\hat{\sigma}_{p}^{2}$ denote variance estimates that will be constructed from a small hold-out set via mom, and define the truncation function $\varphi_{t}(x)=\textup{sgn}(x)(|x|\wedge t)$ for any $x\in\mathbb{R}$ and $t\geq 0$, where $a\wedge b=\min\{a,b\}$. In this notation, the proposed robust max statistic is defined by
@@
With regard to theoretical analysis, we focus on a setting where the tails of the data are quantified by a variant of the $L^{4}$-$L^{2}$ moment equivalence condition, which has gained increasing currency in the high-dimensional robustness literature lugosi2019mean; ke2019user; mendelson2020robust; roy2021empirical; abdalla2022covariance. Specifically, we assume there is some $\delta>0$ such that the bound $\|\langle v,X_{1}-\mu\rangle\|_{L^{4+\delta}}\lesssim\|\langle v,X_{1}-\mu\rangle\|_{L^{2}}$ holds for all $v\in\mathbb{R}^{p}$, and in Proposition 1, we show that this condition is satisfied by heavy-tailed instances of well-known models. The other primary structural assumption in our analysis is that the covariates have a weak variance decay property of the form $\sigma_{(j)}^{2}\asymp j^{-2\beta}$ for some fixed $\beta>0$, where $\sigma_{(1)}^{2}\geq\cdots\geq\sigma_{(p)}^{2}$ are the sorted coordinate-wise variances. Notably, the decay is referred to as weak because the parameter $\beta$ is allowed to be arbitrarily small. Furthermore, it is known that this type of structure arises naturally in a variety of high-dimensional contexts that are related to principal components analysis and functional data analysis, among others (lopes2020bootstrapping). Under the complete set of conditions given in Assumption 1, our main result shows that with high probability, the Kolmogorov distance $\sup_{s\in\mathbb{R}}|\mathbf{P}(\mathcal{M}_{n}\leq s)-\mathbf{P}(\mathcal{M}_{n}^{*}\leq s|X)|$ is nearly of order $n^{-1/2}$, where $\mathbf{P}(\cdot|X)$ denotes probability that is conditional on all of the observations.
From a practical standpoint, the proposed method has several strengths. First, the method does not require fine tuning, which is demonstrated by the fact that we use the simple choices of $\hat{t}_{j}=\sqrt{n}\hat{\sigma}_{j}$ and $\tau=0.9$ throughout all of the experiments presented in Section 4. Second, we show that the method reliably produces well-calibrated tests and confidence intervals across many conditions—including heavy-tailed data generated from separable and elliptical distributions, as well as heavy-tailed functional data with rough sample paths. Third, the simulation results reveal that the proposed method performs favorably in comparison to the pseudomedian approach in (Fan:2023).
@@ 2 Method
Here, we provide the details for constructing the bootstrap sample $\mathcal{M}_{n}^{*}$, as well as simultaneous confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for the coordinate-wise means $\mu_{1},\dots,\mu_{p}$. Further applications of these intervals to various testing problems will be covered later in Section 4.
@@
Next, we turn to the construction of simultaneous confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\mu_{1},\dots,\mu_{p}$. Let $1-\alpha$ denote the nominal simultaneous coverage probability, and let $\hat{q}_{+}(1-\alpha/2)$ denote the empirical $(1-\alpha/2)$-quantile of a collection of bootstrap samples generated in the manner of $\mathcal{M}_{n}^{*}$. Also, let $\underline{\mathcal{M}}_{n}^{*}$ denote the counterpart of $\mathcal{M}_{n}^{*}$ that is obtained by replacing $\max_{1\leq j\leq p}$ with $\min_{1\leq j\leq p}$ in equation (2), and let $\hat{q}_{-}(\alpha/2)$ denote the empirical $(\alpha/2)$-quantile of a collection of bootstrap samples generated in the manner of $\underline{\mathcal{M}}_{n}^{*}$. In this notation, the confidence interval $\hat{\mathcal{I}}_{j}$ is defined by
$\hat{\mathcal{I}}_{j}=\bigg\{x\in\mathbb{R}\,\mathrel{\mathop{\ordinarycolon}}\,\hat{q}_{-}(\alpha/2)\leq\frac{1}{\sqrt{n}\hat{\sigma}_{j}^{\tau}}\sum_{i=1}^{n}\varphi_{\hat{t}_{j}}(X_{ij}-x)\leq\hat{q}_{+}(1-\alpha/2)\bigg\}.$ (7)
Due to the fact that the functions $\varphi_{\hat{t}_{1}}(\cdot),\dots,\varphi_{\hat{t}_{p}}(\cdot)$ are monotone, it is straightforward to compute all the endpoints of $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$.
To comment on the role of the partial standardization parameter $\tau\in[0,1]$, it provides a way to balance two opposing effects that occur in the extreme cases when $\tau$ is equal to 0 or 1. When $\tau=0$, all of the intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ have the same width, which is clearly undesirable when the covariates fluctuate over different scales. Alternatively, when $\tau=1$, all of the covariates will be on approximately “equal footing”, which will tend to make the max statistic $\mathcal{M}_{n}$ sensitive to all $p$ dimensions. This is undesirable in high-dimensional situations where the covariates fluctuate over different scales, because it eliminates a form of low-dimensional structure that can simplify the behavior of $\mathcal{M}_{n}$ when $\tau<1$. To see this, consider a case where $\tau=0$ and $\sigma_{1},\dots,\sigma_{d}$ are much larger than $\sigma_{d+1},\dots,\sigma_{p}$ for some $d\ll p$. In this case, the maximizing index for $\mathcal{M}_{n}$ is likely to reside in the small subset $\{1,\dots,d\}\subset\{1,\dots,p\}$. Thus, the behavior of $\mathcal{M}_{n}$ will be mainly governed by the first $d$ covariates, which intuitively reduces the effective dimension of the problem of approximating the distribution of $\mathcal{M}_{n}$. Accordingly, as was originally proposed in (lopes2020bootstrapping), it is natural to select an intermediate value of $\tau$ between 0 and 1 that can mitigate the unwanted effects that occur at $\tau\in\{0,1\}$.